Diego contacted us after having observed a strange behaviour of the signal TCS_WI_CO2_POWER_CH_PICKOFF measuring the sum of powers of CH, DAS IN and DAS OUT.
A mysterious step at around 11 UTC caused the decrease of the WI CH from 80 mW to ~ 15 mW (see figure). One hypothesis to explain this behaviour is the change in the polarization of the laser but the trigger is not clear and has to be investigated.
We spent 10 min on the bench to verify the proper operation of the power meter and everything was working.
We restored the WI CH nominal power injected into the ITF at 18.50 UTC.